The blocks in a city are all perfect squares. A friend gives you the following directions to a good restaurant; "go north 3 blocks, turn east and go 5 blocks, turn south and go 7 blocks, turn west and go 3 blocks." How far away (straight line distance) is the restaurant?
Answer by CubeyThePenguin(3113)

(note: south is negative of north, west is negative of east)

directions: north 3 blocks, east 5 blocks, north -7 blocks, east -3 blocks
Overall, you are (3-7) = -4 blocks north and (5-3) = 2 blocks east.

distance to start = sqrt((-4)^2 + 2^2) = 2sqrt(5) blocks